OpenStudy (anonymous):

what is the largest natural number that can divides the numbers 1723, 2010, and 5741 and each remainder is 1 ?

OpenStudy (anonymous):

write a computer program for that one

OpenStudy (anonymous):

@ganeshie8 , @waterineyes any idea ? i need a way

OpenStudy (dan815):

factor

OpenStudy (dan815):

u wanna find a natural number that divides all with remainder 1?

OpenStudy (anonymous):

yes divides that numbers and always gives remains 1 respectively

OpenStudy (dan815):

ok well

OpenStudy (dan815):

then just look at umm

OpenStudy (dan815):

1724, 2011, and 5742

OpenStudy (dan815):

factor these out and see largest divisor

OpenStudy (dan815):

wait 2011 is prime

OpenStudy (dan815):

that wud mean ur answer is 1... that cant be it lemme see

OpenStudy (anonymous):

i was represent all number as : 1723 = nk + 1 2010 = np + 1 5741 = nr + 1 with n is divisor, and how to get the largest value of n ?

OpenStudy (dan815):

oh i know lower by 1?

OpenStudy (dan815):

gcd(1722, 2009, and 5740)?

OpenStudy (dan815):

287?

OpenStudy (anonymous):

im not sure --"

OpenStudy (dan815):

ok look

OpenStudy (dan815):

umm

OpenStudy (dan815):

ifff

OpenStudy (dan815):

g=GCD of 1723,2010,5471, which leaves rem 1 then 1723/r = k+1/r 1723= rk+1 1723-1 =rk therefore r must divide 1722 evenly

OpenStudy (dan815):

similarily r will divide all the other numbers -1 evenly

OpenStudy (dan815):

so we simply need to find the GCD of the numbers shifter down by 1

OpenStudy (dan815):

shifted*

OpenStudy (dan815):

GCD(1722, 2009, and 5740)= 287

OpenStudy (dan815):

ok?

OpenStudy (dan815):

where it says g=GCD, it shud say r=GCD

OpenStudy (anonymous):

yes, it is 287.i checked it by calculator... i think i can undestand now, just by using find the GCD thnks, @dan815

OpenStudy (anonymous):

how you can be fast to get GCD(1722, 2009, and 5740)= 287 ?? by using a program comptr ?

OpenStudy (dan815):

umm

OpenStudy (dan815):

ya its not that hard to write a factoring program though especially for such small numbers

OpenStudy (dan815):

sieve of erathonesis

OpenStudy (dan815):

says there mut be a factor that is between 2 to sqrt(1722)

OpenStudy (dan815):

so u check primes between those numbers that divide one u divide then u are checking for 2 to floor sqrt(1722/p1)

OpenStudy (dan815):

primes between there

OpenStudy (dan815):

its pretty fast factoring system

OpenStudy (dan815):

or go to wolfram and type GCD(n1,n2,n3) :P

OpenStudy (dan815):

or just factor like a good boy

OpenStudy (anonymous):

oh, ok... thanks again :) yea, i checked on wolfram: http://www.wolframalpha.com/input/?i=GCD+of+%281722%2C+2009%2C+and+5740%29&dataset=&equal=Submit i dont like do that like a good boy ;P

OpenStudy (dan815):

1722=2*861 =2*3*287

OpenStudy (dan815):

8+6+1 = 15 therefore 3 must factor

OpenStudy (dan815):

if have n1*100+n2*10+n3*1 then (n1*100+n2*10+n3*1)/3 = n1*(33+1/3)+n2(3+1/3)+n3/3 =33n1+3n2+n3/3+n1/3+n2/3 =33n1+3n2+ (n1+n2+n3)/3 therefore as long as n1+n2+n3 add upto a multiple of 3

OpenStudy (dan815):

u can see how this same pattern follows infinite digits

ganeshie8 (ganeshie8):

1723 = nk + 1 2010 = np + 1 5741 = nr + 1 n | (1723-1) n | (2010-1) n | (5741-1)

OpenStudy (anonymous):

yes, i caught that @ganeshie8 . thanks :)

ganeshie8 (ganeshie8):

euclid algorithm is superior to prime factorization for finding gcd

OpenStudy (dan815):

what is that algoritm

ganeshie8 (ganeshie8):

2009 = 1722*1 + 287 1722 = 287*6 + 0 so gcd(2009,1722) = 287

OpenStudy (dan815):

:O coool

ganeshie8 (ganeshie8):

since 287 | 5743, we are done.